You find that the annual standard deviation of a stock's returns is equal to 35%. For a 4 year holding period the standard deviation of your total return would equal ________.
Answer #1

Standard deviation for 4 year=sqrt(4)*standard deviation of 1 year=2*standard deviation opf 1 year=2*35%=70%